Show steps on how you got the answers:

1. Busy Bee Industries installed a machine in its factory at a cost of $84,000 on May 1, 2017. The machine useful life is estimated at eight years with a $9,000 salvage value. Determine the machine's 2017 depreciation using the double-declining-balance method of depreciation.

a) $14,000

b) $10,500

c) $21,000

d) 12,500

e) 18,750

2. Spacely's Sprockets purchased a machine on September 1 for $400,000. The machine's useful life was estimated at six years or 500,000 units of the product with a $25,000 trade-in value. During its second year, the machine produced 87,000 units of product. Assuming units-of-production depreciation, calculate the machine's second-year depreciation.

a) $69,600

b) 66,667

c) 65,250

d) $62,500

e) $31,250

3. Cherokee Company sold a bulldozer that originally cost $16,000 for $14,200. Accumulated depreciation on this bulldozer was $2,000. The gain or loss from the sale is:

a) No gain or loss

b) $200 gain

c) $200 loss

d) $1,800 loss

e) $16,000 loss

4. Giorgio Company estimates that future costs to satisfy its product warranty obligation amount to 3% of sales. In January the company sold merchandise for $50,000 cash and paid $1,200 to repair products returned for warranty work. Prepare the journal entries to record (1) the accrual of warranty costa and (2) the repair of products for warranty work.
Answer #1

1. Double declining rate = 200/useful life = 200/8 = 25%

Depreciation = 84,000*25%*8/12

= 14,000

Option A

2. Depreciation = (Cost - Salvage value) /life in units * Units produced

= (400,000-25,000)/500,000 * 87,000

= 0.75*87,000

= 65,250

Option C

3. Gain(Loss) = Cash received - Book value

= 14,200 - (16,000-2000)

= 14,200 - 14,000

= 200 gain

Option B